# Harnessing Diffusion-Generated Synthetic Images for Fair Image Classification

### v1

@@ 1 Introduction
Image classification models often exhibit harmful biases, posing significant risks for real-world deployment (wang2021gender; zhao2017men; metaxa2021image). These biases arise from imbalances in training data; e.g., in CelebA (celeba), blond female faces considerably outnumber blond males, leading to misclassification of the latter. While numerous debiasing techniques have been proposed (sagawa2019distributionally; kirichenko2022last; nam2020learning), mitigating bias becomes increasingly tough when dataset imbalances become severe. With the recent breakthroughs in image generation using models like Stable Diffusion (ramesh2022hierarchical), we pose a critical question: Can we harness the generative power of such models to create images that facilitate the training of fair classification systems, even in presence of extreme dataset bias?
A recent work FFR (qraitem2023fake) leverages images generated by Stable Diffusion (SD) to train fair classification systems. However, due to the stochastic nature of diffusion models (shin2023fill), SD-generated images often diverge from the original data distribution and may not follow prompt instructions accurately. Consider the Waterbirds dataset (sagawa2019distributionally), which contains ‘waterbird’ and ‘landbird’ classes. Spurious correlations arise as classification models associate water backgrounds with waterbirds and land backgrounds with landbirds, thus relying on background rather than the bird features for predictions. Attempts to prompt SD to generate ‘waterbird on land’ images often produce water backgrounds, even when explicitly instructed otherwise (Appendix Fig. 8). FFR attempts to mitigate this by using highly specific prompts, such as ‘photo of a flamingo on pavement’, relying on exact bird names and background details. However, without precise domain knowledge, such prompts may yield irrelevant or out-of-distribution images. In this work, we explore fine-tuning generative models directly on the dataset, enabling them to better capture the data distribution and significantly improve classification fairness.
@@
To generate images closely aligned with the training set, we fine-tune Stable Diffusion using LoRA (lora) on each dataset group (e.g., Non-Blond Females, Blond Males in CelebA (celeba)). Prior work on synthetic data augmentation (shin2023fill) trains personalized diffusion models on individual classes to produce class-consistent, in-distribution images. Building on this, we further explore DreamBooth (ruiz2022dreambooth), which introduces a special token (‘[V]’) to represent the subject, improving resemblance to real samples. While such models typically target visually similar objects (e.g., a specific dog breed), our training groups contain diverse images sharing only a common attribute (e.g., hair color). To address this intra-group variation, we explore a simple extension—Clustered DreamBooth—which clusters each group into visually similar subsets and fine-tunes a separate DreamBooth model per cluster. These different generation methods are illustrated in Fig. 1. Using each strategy, we generate equal number of images per group.
After generating group-wise images, the key question becomes: how should we use them for training? FFR addresses this by first training a classification model on the synthetic data, followed by finetuning on the original dataset. However, since the original data is biased, this step risks reintroducing bias—unless extensive hyperparameter tuning is performed, as in FFR. To avoid this, we propose to finetune only the classifier’s softmax layer with the real data (basu2024mitigating). Both the training and finetuning stages are optimized using a weighted combination of Cross-Entropy and Supervised Contrastive Loss (khosla2020supervised), which enhances feature representations by promoting better class separation.
We evaluate our proposed pipeline on three fairness benchmarks—Waterbirds (sagawa2019distributionally), CelebA (celeba), and UTKFace (zhang2017age). The training on vanilla SD generated images followed by final layer retraining improves the accuracy of the classification systems over FFR indicating the utility of our finetuning approach. The different finetuning strategies discussed in the paper also offer added advantage for all the datasets, performing comparably to traditional debiasing techniques like GroupDRO (sagawa2019distributionally) and SELF (labonte2023towards). Notably, as dataset biases intensify, such methods deteriorate, while our generation-based approaches remain robust, demonstrating their effectiveness in highly imbalanced settings. We summarize the key contributions below:
@@ 2 Related Work
Bias Mitigation has been widely studied, with approaches falling into two categories: known and unknown biases. For known biases, the spurious attribute is known apriori (kim2019learning; li2019repair; sagawa2019distributionally; arjovsky2019invariant; teney2021unshuffling; tartaglione2021end; wang2020towards; Wang_2022_CVPR; basu2023rmlvqa). In GroupDRO (sagawa2019distributionally), the worst-group training loss is optimized. Last Layer Retraining (kirichenko2022last) shows that pretraining the model on the biased dataset and then only retraining the final classification layer with a group-balanced validation set can help debias the model. Semi-supervised approaches assume bias annotations only for a few samples (nam2020learning; jung2022learning). For unknown biases, where bias attributes and labels are not available (creager2021environment; lee2021learning; Li_2022_ECCV; lahoti2020fairness; ahn2023mitigating; liu2021just; huang2020self; hong2021unbiased; darlow2020latent; basu2024mitigating), some approaches use dual-branch networks—one amplifying bias, the other mitigating it (nam2020learning; lee2021learning; liu2022avoiding). Contrastive-based methods (zhang2022correct; zhang2022contrastive) refine feature representations by clustering same-class samples, identifying pseudo bias labels via model misclassifications. Such biases are also exhibited by diffusion models (basu2023inspecting; parihar2024balancing).
@@
Data Augmentation using Generative Models. Many recent works utilize generative models for data augmentation (trabucco2023effective; azizi2023synthetic; du2024dream; zheng2023toward; mariani2018bagan). The trend began with GANs (goodfellow2020generative). BAGAN (mariani2018bagan) was used to augment class-imbalanced datasets to enhance minority class performance. With diffusion models, DA-Fusion (trabucco2023effective) employs Textual Inversion (gal2022image) to generate augmentations for images of every class, and then during training, in each batch, retain every original image with probability $p$ and an augmented image with probability (1p)(1-p). DiffuseMix (Islam_2024_CVPR) combines a partial natural image and its generated counterpart from the diffusion model, and thereafter combats adversarial attacks by blending a randomly selected structural pattern from a set of fractal images into the concatenated image to form the final augmented version for training.
@@
Generative Models for Debiasing utilize generative models to debias classification systems (an20222; ramaswamy2021fair; qraitem2023fake; sharmanska2020contrastive). GAN-based approaches (an20222; ramaswamy2021fair) train generative models on the training images to synthesize bias-conflicting samples that can augment the original data. The diffusion-based methods do not train the models from scratch, rather manipulate existing pretrained models to generate group-balanced images. FFR (qraitem2023fake) likewise generates group-balanced images from Stable Diffusion, trains the classification model on this synthetic data, before finetuning the latter with real data. Our work additionally explores the use of different finetuning mechanisms in text-to-image models to generate in-distribution images directly from the training image groups.
@@ 3.1 Generating Synthetic Images
With advancements in generative modeling (song2020denoising; ramesh2022hierarchical), we explore their potential in training fair classifiers by generating images that reflect the training distribution and improve generalization to minority groups. Below, we outline our image generation methods and strategies for leveraging synthetic data to enhance fairness.
@@
Dreambooth. To strengthen the resemblance between the training and generated images, we explore Dreambooth (ruiz2022dreambooth), a text-to-image personalization model that learns to imitate an object or a concept (e.g., a specific dog) from a small set of images depicting that object. It finetunes a pretrained text-to-image model by learning a unique identifier (e.g., “[V]”) such that on inference time, if the model is queried by that identifier (e.g., “photo of a [V] dog”), it generates new images of the given object. Likewise, we sample $100$ images from each training group, and train a separate Dreambooth model $h$ on each group, where the prompt is of the form “photo of a [V] y”.
Clustered Dreambooth. Dreambooth specializes in learning a concept from 353-5 images. However, a training group like Blond Male consists of images of many individuals sharing a common trait, hair color. To prevent overwhelming a single Dreambooth model with multiple images of a training group, we explore a simple extension of first clustering the CLIP embeddings (pmlr-v139-radford21a) of the images within each group. Let $k^{g}_{D}$ denote the number of clusters, where $D$ and $g$ refer to the training dataset and a group in $D$ respectively. We train a pool of Dreambooth models $\mathcal{H}^{g}=\{h^{g}_{1},h^{g}_{2},h^{g}_{3},\cdots,h^{g}_{k^{g}_{D}}\}$ on the obtained clusters. We implement Clustered-Dreambooth (i.e., the Dreambooth pool $\mathcal{H}^{g}$) using LoRA-based finetuning (hu2021lora), which ensures lesser, feasible training time. Finally the trained models are utilized to generate images for each $g$. For simplicity, we assume equal $k^{g}_{D}$ for each group $g$, and denote the number of clusters as $k_{D}$ for the rest of the paper.
@@ 3.3 Stage 2: Finetuning with Original Data
We train both stages using a weighted sum of CE loss and Supervised Contrastive (SupCon) loss (khosla2020supervised) (see Appendix A.1) to enhance class separation in both stages: $L=\beta.L_{\text{CE}}+(1-\beta)L_{\text{sup-con}}$, where $\beta=0.5$ in all experiments.
@@ 4 Experiments and Results
Datasets For Evaluation: Waterbirds (sagawa2019distributionally) is a dataset of bird images, labeled as waterbird if the bird is commonly seen with waterbodies on background, and landbird otherwise. The dataset, with $4,795$ samples, suffers from background bias: only a few waterbirds in the dataset have land on the background, and few landbirds have water. We also consider two real-world datasets. The CelebA dataset (celeba) consists of $202,599$ images of celebrities with annotations of $40$ binary attributes. We choose Blond Hair as the target attribute, which is known to suffer from gender biases (Li_2022_ECCV; seo2022unsupervised). UTKFace (zhang2017age) is a dataset of human faces, having $20,000$ images with annotations of age, gender, and ethnicity. We use gender and age as the target and bias attributes respectively – where female adults dominate female children, and male children dominate male adults. Further details on this data split is provided in Appendix A.2. For each dataset, we report the worst (WGA) and average group (AGA) accuracies, following previous works (sagawa2019distributionally; labonte2023towards).
@@
Dreambooth (ruiz2022dreambooth) $82.1^{\pm 0.00}$ $\underline{87.9}^{\pm 0.32}$ $\underline{78.8}^{\pm 0.21}$ $84.6^{\pm 0.19}$ $\underline{80.4}$ $\underline{86.2}$
@@
Dreambooth (ruiz2022dreambooth) $57.9^{\pm 3.91}$ $80.9^{\pm 0.76}$ $\underline{66.9}^{\pm 2.59}$ $77.1^{\pm 0.26}$ $62.4$ $79.0$
@@
Implementation Details: We use SD v1.4 for all experiments. The prompts used to generate the images for each variant is provided in Appendix A.3. For each dataset group, we generate $M=5000$ images for each method, and score the images using the CLIP-Score defined in eq. 3.2. For simplicity, we set the weighting parameter $\alpha$ to be $0.5$ for the original dataset ($\alpha=1$ for the severely biased version). Accordingly, we select the top $75\%$ images from each group. An ImageNet (deng2009imagenet)-pretrained ResNet-50 (He_2016_CVPR) model is trained on the group-balanced data with CE and SupCon losses. The model is trained for 20 epochs using the SGD optimizer during both the training and finetuning stages. As the validation sets of the datasets are not group-balanced, we refrain from experimenting with hyperparameters, and uniformly set a learning rate of $1\text{e}-3$, weight decay of $1\text{e}-3$ and batch size of $128$. We discuss the choice of clusters for the Clustered Dreambooth method in § 4.2.
@@ 4.1 Quantitative Results
Our goal is to obtain rich representations from synthetic images to enable classification models to be fair, even when finetuned on biased real data. Table 1 presents the worst and average group accuracies (WGA and AGA) across all test groups and dataset variants. To further motivate the importance of the generative models in training fair classifiers, we compare their performance with traditional debiasing methods like Group-DRO (GDRO) (sagawa2019distributionally) and SELF (labonte2023towards). Key findings are summarized below:
Performance on the Original Dataset. Our classification pipeline, leveraging images from vanilla SD, outperforms FFR (qraitem2023fake) by an average of $2.4\%$. This improvement can be attributed to our last layer retraining-based finetuning stage. DreamBooth achieves the highest worst-group accuracy for Waterbirds ($89.3\%$), followed by Clustered DreamBooth ($88.1\%$) and LoRA fine-tuning ($86.5\%$). For CelebA and UTKFace, Clustered DreamBooth outperforms all other generative methods and is the only one to surpass ERM scores for UTKFace, improving upon LoRA by $7.4\%$. This highlights the advantage of training multiple DreamBooth models on clustered subsets, particularly for facial datasets. While Clustered DreamBooth has higher time complexity, $k_{D}$ can be adjusted based on resource constraints, with $k_{D}=1$ (a single DreamBooth model per group) as a special case. Notably, since DreamBooth models are themselves trained with LoRA, per-model training remains computationally efficient. Overall, fine-tuning consistently outperforms vanilla Stable Diffusion across all datasets, achieving accuracy comparable to traditional debiasing methods like Group-DRO.
@@
Generative-Based Pretraining + GDRO Finetuning ($\mathbf{0.999}$ bias ratio). Table 1 shows that the WGA of GDRO (sagawa2019distributionally) suffers for high bias ratio. We next analyse the impact of finetuning the classification model pretrained by the Clustered Dreambooth images  [^1] We choose Clustered Dreambooth for different ablations and experiments only as an example. for Waterbirds and CelebA, using the GDRO loss, leading to a $58.3\%$ and $37.2\%$ performance increase for Waterbirds and CelebA respectively (Table 2). This analysis further highlights the positive effect of group-balanced synthetic image pretraining. Qualitative examples are shown in Appendix Fig. 10.
@@
Time Complexity of Finetuning. We analyze the time complexity vs performance tradeoff for the finetuning techniques, averaged across all datasets and bias ratios. Clustered DreamBooth outperforms others, but it has higher time complexity of $\mathcal{O}(|\mathcal{G}_{D}|.k_{D})$, where $|\mathcal{G}_{D}|$ represents the number of groups in the training set, and $k_{D}:$ the number of clusters per group. In contrast, vanilla SD, which requires no finetuning, yields the lowest score across all datasets. These tradeoffs are presented in Table 3, providing insights to help practitioners select the most suitable method based on their application needs. We also evaluate distribution similarities using Fréchet Inception Distance (FID) (heusel2017gans) between generated and real CelebA images: finetuning approaches outperform FFR significantly (see Appendix Table 11).

### v2

@@ 1 Introduction
Image classification models often exhibit harmful biases, posing significant risks for real-world deployment (Wang, Liu, and Wang 2021; Zhao et al. 2017; Metaxa et al. 2021). These biases arise from imbalances in training data; e.g., in CelebA (Celeba ), blond female faces considerably outnumber blond males, leading to misclassification of the latter. While numerous debiasing techniques have been proposed (Sagawa et al. 2019; Kirichenko, Izmailov, and Wilson 2022; Nam et al. 2020), mitigating bias becomes increasingly tough when dataset imbalances become severe. With the recent breakthroughs in image generation using models like Stable Diffusion (Ramesh et al. 2022), we pose a critical question: Can we harness the generative power of such models to create images that facilitate the training of fair classification systems, even in presence of extreme dataset bias?
A recent work FFR (Qraitem, Saenko, and Plummer 2023b) leverages images generated by Stable Diffusion (SD) to train fair classification systems. However, due to the stochastic nature of diffusion models (Shin, Kang, and Park 2023), SD-generated images often diverge from the original data distribution and may not follow prompt instructions accurately. Consider the Waterbirds dataset (Sagawa et al. 2019), which contains ‘waterbird’ and ‘landbird’ classes. Spurious correlations arise as classification models associate water backgrounds with waterbirds and land backgrounds with landbirds, thus relying on background rather than the bird features for predictions. Attempts to prompt SD to generate ‘waterbird on land’ images often produce water backgrounds, even when explicitly instructed otherwise (Appendix Fig. 8). FFR attempts to mitigate this by using highly specific prompts, such as ‘photo of a flamingo on pavement’, relying on exact bird names and background details. However, without precise domain knowledge, such prompts may yield irrelevant or out-of-distribution images. In this work, we explore fine-tuning generative models directly on the dataset, enabling them to better capture the data distribution and significantly improve classification fairness.
@@
To generate images closely aligned with the training set, we fine-tune Stable Diffusion using LoRA (Hu et al. 2021a) on each dataset group (e.g., Non-Blond Females, Blond Males in CelebA (Celeba )). Prior work on synthetic data augmentation (Shin, Kang, and Park 2023) trains personalized diffusion models on individual classes to produce class-consistent, in-distribution images. Building on this, we further explore DreamBooth (Ruiz et al. 2022), which introduces a special token (‘[V]’) to represent the subject, improving resemblance to real samples. While such models typically target visually similar objects (e.g., a specific dog breed), our training groups contain diverse images sharing only a common attribute (e.g., hair color). To address this intra-group variation, we explore a simple extension—Clustered DreamBooth—which clusters each group into visually similar subsets and fine-tunes a separate DreamBooth model per cluster. These different generation methods are illustrated in Fig. 1. Using each strategy, we generate equal number of images per group.
After generating group-wise images, the key question becomes: how should we use them for training? FFR addresses this by first training a classification model on the synthetic data, followed by finetuning on the original dataset. However, since the original data is biased, this step risks reintroducing bias—unless extensive hyperparameter tuning is performed, as in FFR. To avoid this, we propose to finetune only the classifier’s softmax layer with the real data (Basu, Mallick et al. 2024). Both the training and finetuning stages are optimized using a weighted combination of Cross-Entropy and Supervised Contrastive Loss (Khosla et al. 2020), which enhances feature representations by promoting better class separation.
We evaluate our proposed pipeline on three fairness benchmarks—Waterbirds (Sagawa et al. 2019), CelebA (Celeba ), and UTKFace (Zhang, Song, and Qi 2017). The training on vanilla SD generated images followed by final layer retraining improves the accuracy of the classification systems over FFR indicating the utility of our finetuning approach. The different finetuning strategies discussed in the paper also offer added advantage for all the datasets, performing comparably to traditional debiasing techniques like GroupDRO (Sagawa et al. 2019) and SELF (LaBonte, Muthukumar, and Kumar 2023). Notably, as dataset biases intensify, such methods deteriorate, while our generation-based approaches remain robust, demonstrating their effectiveness in highly imbalanced settings. We summarize the key contributions below:
@@ 2 Related Work
Bias Mitigation has been widely studied, with approaches falling into two categories: known and unknown biases. For known biases, the spurious attribute is known apriori (Kim et al. 2019; Li and Vasconcelos 2019; Sagawa et al. 2019; Arjovsky et al. 2019; Teney, Abbasnejad, and van den Hengel 2021; Tartaglione, Barbano, and Grangetto 2021; Wang et al. 2020, 2022; Basu, Addepalli, and Babu 2023). In GroupDRO (Sagawa et al. 2019), the worst-group training loss is optimized. Last Layer Retraining (Kirichenko, Izmailov, and Wilson 2022) shows that pretraining the model on the biased dataset and then only retraining the final classification layer with a group-balanced validation set can help debias the model. Semi-supervised approaches assume bias annotations only for a few samples (Nam et al. 2020; Jung, Chun, and Moon 2022). For unknown biases, where bias attributes and labels are not available (Creager, Jacobsen, and Zemel 2021; Lee et al. 2021; Li, Hoogs, and Xu 2022; Lahoti et al. 2020; Ahn, Kim, and Yun 2023; Liu et al. 2021; Huang et al. 2020; Hong and Yang 2021; et al. 2020; Basu, Mallick et al. 2024), some approaches use dual-branch networks—one amplifying bias, the other mitigating it (Nam et al. 2020; Lee et al. 2021; Liu et al. 2022). Contrastive-based methods (Zhang et al. 2022; Zhang and Ré 2022) refine feature representations by clustering same-class samples, identifying pseudo bias labels via model misclassifications. Such biases are also exhibited by diffusion models (Basu, Babu, and Pruthi 2023; Parihar et al. 2024).
@@
Data Augmentation using Generative Models. Many recent works utilize generative models for data augmentation (Trabucco et al. 2023; Azizi et al. 2023; Du et al. 2024; Zheng, Wu, and Li 2023; Mariani et al. 2018). The trend began with GANs (Goodfellow et al. 2020). BAGAN (Mariani et al. 2018) was used to augment class-imbalanced datasets to enhance minority class performance. With diffusion models, DA-Fusion (Trabucco et al. 2023) employs Textual Inversion (Gal et al. 2022) to generate augmentations for images of every class, and then during training, in each batch, retain every original image with probability $p$ and an augmented image with probability (1p)(1-p). DiffuseMix (Islam et al. 2024) combines a partial natural image and its generated counterpart from the diffusion model, and thereafter combats adversarial attacks by blending a randomly selected structural pattern from a set of fractal images into the concatenated image to form the final augmented version for training.
@@
Generative Models for Debiasing utilize generative models to debias classification systems (An et al. 2022; Ramaswamy, Kim, and Russakovsky 2021; Qraitem, Saenko, and Plummer 2023b; Sharmanska et al. 2020). GAN-based approaches (An et al. 2022; Ramaswamy, Kim, and Russakovsky 2021) train generative models on the training images to synthesize bias-conflicting samples that can augment the original data. The diffusion-based methods do not train the models from scratch, rather manipulate existing pretrained models to generate group-balanced images. FFR (Qraitem, Saenko, and Plummer 2023b) likewise generates group-balanced images from Stable Diffusion, trains the classification model on this synthetic data, before finetuning the latter with real data. Our work additionally explores the use of different finetuning mechanisms in text-to-image models to generate in-distribution images directly from the training image groups.
@@ 3.1 Generating Synthetic Images
With advancements in generative modeling (Song, Meng, and Ermon 2020; Ramesh et al. 2022), we explore their potential in training fair classifiers by generating images that reflect the training distribution and improve generalization to minority groups. Below, we outline our image generation methods and strategies for leveraging synthetic data to enhance fairness.
@@
Dreambooth. To strengthen the resemblance between the training and generated images, we explore Dreambooth (Ruiz et al. 2022), a text-to-image personalization model that learns to imitate an object or a concept (e.g., a specific dog) from a small set of images depicting that object. It finetunes a pretrained text-to-image model by learning a unique identifier (e.g., “[V]”) such that on inference time, if the model is queried by that identifier (e.g., “photo of a [V] dog”), it generates new images of the given object. Likewise, we sample $100$ images from each training group, and train a separate Dreambooth model $h$ on each group, where the prompt is of the form “photo of a [V] y”.
Clustered Dreambooth. Dreambooth specializes in learning a concept from 353-5 images. However, a training group like Blond Male consists of images of many individuals sharing a common trait, hair color. To prevent overwhelming a single Dreambooth model with multiple images of a training group, we explore a simple extension of first clustering the CLIP embeddings (Radford et al. 2021) of the images within each group. Let $k^{g}_{D}$ denote the number of clusters, where $D$ and $g$ refer to the training dataset and a group in $D$ respectively. We train a pool of Dreambooth models $\mathcal{H}^{g}=\{h^{g}_{1},h^{g}_{2},h^{g}_{3},\cdots,h^{g}_{k^{g}_{D}}\}$ on the obtained clusters. We implement Clustered-Dreambooth (i.e., the Dreambooth pool $\mathcal{H}^{g}$) using LoRA-based finetuning (Hu et al. 2021b), which ensures lesser, feasible training time. Finally the trained models are utilized to generate images for each $g$. For simplicity, we assume equal $k^{g}_{D}$ for each group $g$, and denote the number of clusters as $k_{D}$ for the rest of the paper.
@@ 3.3 Stage 2: Finetuning with Original Data
We train both stages using a weighted sum of CE loss and Supervised Contrastive (SupCon) loss (Khosla et al. 2020) (see Appendix A.1) to enhance class separation in both stages: $L=\beta.L_{\text{CE}}+(1-\beta)L_{\text{sup-con}}$, where $\beta=0.5$ in all experiments.
@@ 4 Experiments and Results
Datasets For Evaluation: Waterbirds (Sagawa et al. 2019) is a dataset of bird images, labeled as waterbird if the bird is commonly seen with waterbodies on background, and landbird otherwise. The dataset, with $4,795$ samples, suffers from background bias: only a few waterbirds in the dataset have land on the background, and few landbirds have water. We also consider two real-world datasets. The CelebA dataset (Celeba ) consists of $202,599$ images of celebrities with annotations of $40$ binary attributes. We choose Blond Hair as the target attribute, which is known to suffer from gender biases (Li, Hoogs, and Xu 2022; Seo, Lee, and Han 2022). UTKFace (Zhang, Song, and Qi 2017) is a dataset of human faces, having $20,000$ images with annotations of age, gender, and ethnicity. We use gender and age as the target and bias attributes respectively – where female adults dominate female children, and male children dominate male adults. Further details on this data split is provided in Appendix A.2. For each dataset, we report the worst (WGA) and average group (AGA) accuracies, following previous works (Sagawa et al. 2019; LaBonte, Muthukumar, and Kumar 2023).
@@
Dreambooth (Ruiz et al. 2022) $82.1^{\pm 0.00}$ $\underline{87.9}^{\pm 0.32}$ $\underline{78.8}^{\pm 0.21}$ $84.6^{\pm 0.19}$ $\underline{80.4}$ $\underline{86.2}$
@@
Dreambooth (Ruiz et al. 2022) $57.9^{\pm 3.91}$ $80.9^{\pm 0.76}$ $\underline{66.9}^{\pm 2.59}$ $77.1^{\pm 0.26}$ $62.4$ $79.0$
@@
Implementation Details: We use SD v1.4 for all experiments. The prompts used to generate the images for each variant is provided in Appendix A.3. For each dataset group, we generate $M=5000$ images for each method, and score the images using the CLIP-Score defined in eq. 3.2. For simplicity, we set the weighting parameter $\alpha$ to be $0.5$ for the original dataset ($\alpha=1$ for the severely biased version). Accordingly, we select the top $75\%$ images from each group. An ImageNet (Deng et al. 2009)-pretrained ResNet-50 (He et al. 2016) model is trained on the group-balanced data with CE and SupCon losses. The model is trained for 20 epochs using the SGD optimizer during both the training and finetuning stages. As the validation sets of the datasets are not group-balanced, we refrain from experimenting with hyperparameters, and uniformly set a learning rate of $1\text{e}-3$, weight decay of $1\text{e}-3$ and batch size of $128$. We discuss the choice of clusters for the Clustered Dreambooth method in § 4.2.
@@ 4.1 Quantitative Results
Our goal is to obtain rich representations from synthetic images to enable classification models to be fair, even when finetuned on biased real data. Table 1 presents the worst and average group accuracies (WGA and AGA) across all test groups and dataset variants. To further motivate the importance of the generative models in training fair classifiers, we compare their performance with traditional debiasing methods like Group-DRO (GDRO) (Sagawa et al. 2019) and SELF (LaBonte, Muthukumar, and Kumar 2023). Key findings are summarized below:
Performance on the Original Dataset. Our classification pipeline, leveraging images from vanilla SD, outperforms FFR (Qraitem, Saenko, and Plummer 2023b) by an average of $2.4\%$. This improvement can be attributed to our last layer retraining-based finetuning stage. DreamBooth achieves the highest worst-group accuracy for Waterbirds ($89.3\%$), followed by Clustered DreamBooth ($88.1\%$) and LoRA fine-tuning ($86.5\%$). For CelebA and UTKFace, Clustered DreamBooth outperforms all other generative methods and is the only one to surpass ERM scores for UTKFace, improving upon LoRA by $7.4\%$. This highlights the advantage of training multiple DreamBooth models on clustered subsets, particularly for facial datasets. While Clustered DreamBooth has higher time complexity, $k_{D}$ can be adjusted based on resource constraints, with $k_{D}=1$ (a single DreamBooth model per group) as a special case. Notably, since DreamBooth models are themselves trained with LoRA, per-model training remains computationally efficient. Overall, fine-tuning consistently outperforms vanilla Stable Diffusion across all datasets, achieving accuracy comparable to traditional debiasing methods like Group-DRO.
@@
Generative-Based Pretraining + GDRO Finetuning ($\mathbf{0.999}$ bias ratio). Table 1 shows that the WGA of GDRO (Sagawa et al. 2019) suffers for high bias ratio. We next analyse the impact of finetuning the classification model pretrained by the Clustered Dreambooth images  [^1] We choose Clustered Dreambooth for different ablations and experiments only as an example. for Waterbirds and CelebA, using the GDRO loss, leading to a $58.3\%$ and $37.2\%$ performance increase for Waterbirds and CelebA respectively (Table 2). This analysis further highlights the positive effect of group-balanced synthetic image pretraining. Qualitative examples are shown in Appendix Fig. 10.
@@
Time Complexity of Finetuning. We analyze the time complexity vs performance tradeoff for the finetuning techniques, averaged across all datasets and bias ratios. Clustered DreamBooth outperforms others, but it has higher time complexity of $\mathcal{O}(|\mathcal{G}_{D}|.k_{D})$, where $|\mathcal{G}_{D}|$ represents the number of groups in the training set, and $k_{D}:$ the number of clusters per group. In contrast, vanilla SD, which requires no finetuning, yields the lowest score across all datasets. These tradeoffs are presented in Table 3, providing insights to help practitioners select the most suitable method based on their application needs. We also evaluate distribution similarities using Fréchet Inception Distance (FID) (Heusel et al. 2017) between generated and real CelebA images: finetuning approaches outperform FFR significantly (see Appendix Table 11).